OpenStudy (anonymous):

@iGreen. A cylindrical cardboard mailing tube has a diameter of 3 in. and a height of 28 in. A total of 23% of the tube’s capacity is filled with spherical packing foam. Each foam sphere has a diameter of 0.8 in. Approximately how many pieces of foam are in the tube? Use 3.14 to approximate pi and round your answer to the nearest whole piece of foam. ___

OpenStudy (anonymous):

Hmm..let's start by finding the volume of the mailing tube. \(V = \pi r^2 h\) \(V = (3.14)(1.5^2)(28)\) Simplify

OpenStudy (anonymous):

3.14 times 2.25 times 28

OpenStudy (anonymous):

197.82

OpenStudy (anonymous):

Yes, and 23% of 197.82 is 45.4986. So 45.4986 of the volume is packing foam. Now let's find the volume of each piece of foam. \(V = \dfrac{4}{3} \pi r^2\) \(V = \dfrac{4}{3} (3.14)(0.4^2)\)

OpenStudy (anonymous):

1.3333 times 3.14 times 0.16

OpenStudy (anonymous):

0.66984992

OpenStudy (anonymous):

Yes, now divide: 45.4986 / 0.66984992

OpenStudy (anonymous):

@Master_Chief_117

OpenStudy (anonymous):

(Sorry. nephew yelling in ear)

OpenStudy (anonymous):

cant really think. 67.9235730893

OpenStudy (anonymous):

Yep, so it should be 68.

OpenStudy (anonymous):

The answer? (Sorry...still dealing with a yelling neph)

OpenStudy (anonymous):

I just said 68 :P

OpenStudy (anonymous):

You said 68 was the answer?

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

Thanks. i have bad net. this might be it for the day. Thanks again

OpenStudy (anonymous):

i dont think thats right

OpenStudy (moses238):

they did v = 4/3 * 3.14 * 0.4^2 they were needed to do v = 4/3 * 3.14 * 0.04^3